OpenStudy (anonymous):

What is the length of the diameter of the base of a cylinder if its volume is 768Pi cubic inches and its height is 12 inches

jimthompson5910 (jim_thompson5910):

V = pi*r^2*h 768pi = pi*r^2*12 solve for r, then use the value of r to find d = 2r

OpenStudy (anonymous):

I dont know how to solve for r

jimthompson5910 (jim_thompson5910):

768pi = pi*r^2*12 768pi/pi = pi*r^2*12/pi 768 = 12r^2 12r^2 = 768 how about now?

OpenStudy (anonymous):

12r/12=27.7/12 r=2.3 Am i right?

jimthompson5910 (jim_thompson5910):

12r^2 = 768 r^2 = 768/12 r^2 = 64 r = ???

OpenStudy (anonymous):

r=8 !

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

so the answer is 8?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

oh wait

OpenStudy (anonymous):

Thankssss

jimthompson5910 (jim_thompson5910):

the radius is 8

jimthompson5910 (jim_thompson5910):

the diameter isn't 8 though

OpenStudy (anonymous):

oh, the diameter is 16. d=2r

jimthompson5910 (jim_thompson5910):

yep you got it